OpenStudy (anonymous):

Rectangle A has an area of 16 - x2. Rectangle B has an area of x2 + 2x - 24. In simplest form, what is the ratio of the area of Rectangle A to the area of Rectangle B? Show your work.

OpenStudy (anonymous):

@wolf1728

OpenStudy (wolf1728):

Here I am

OpenStudy (anonymous):

Loll welcome

OpenStudy (wolf1728):

16 - x2 = (4 +x) • (4 -x) x2 + 2x - 24 = (x +6) • (x -4) That's a start

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

dont we have to cross out the 4

OpenStudy (anonymous):

no forget it

OpenStudy (wolf1728):

Should I divide 1 by the other? (x +6) • (x -4) / (4 +x) • (4 -x)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

= (-x - 4)/(x + 6

OpenStudy (wolf1728):

Seems that nothing cancels out in that division does it?

OpenStudy (anonymous):

no

OpenStudy (wolf1728):

Maybe there's another way

OpenStudy (anonymous):

sorry serious lag

OpenStudy (wolf1728):

If we get the roots of the equations x= -6 and x = 4 Does that help?

OpenStudy (anonymous):

i'm gonna be honest im am not sure

OpenStudy (wolf1728):

maybe there's more I can do

OpenStudy (anonymous):

maybe..

OpenStudy (anonymous):

well you have done a lot its just me i'm beat

OpenStudy (wolf1728):

Well, I was thinking what might be the factors of an area of 16-x² The root of that equation is 4 and so Area = 16 -4² = 0 The other root -4 makes the Area = 16 -(-4²) = 0 Something seems strange about that 16 -x² equation

OpenStudy (anonymous):

what seems strange?

OpenStudy (wolf1728):

How about x² + 2x - 24 = 16 - x² That can't be correct because the 2 areas are different. ******************************************************* What seems strange about 16 -x² is that the roots of it make the Area = 0

OpenStudy (anonymous):

Had a feeling okay so how do we make this right?

OpenStudy (wolf1728):

I don't know. How about the area of the rectangle that is x² + 2x - 24 The one positive root of that is 4 so let's say Area = 4² +2 • 4 + 24 Area = 16 + 8 + 24 Area = 48 That seems correct

OpenStudy (anonymous):

Hmm it does seem correct

OpenStudy (wolf1728):

Nope - there's a mistake Area = 4² +2 • 4 + 24 is wrong SHOULD BE -24

OpenStudy (wolf1728):

That makes the other area = 0 also. This is a weird problem

OpenStudy (anonymous):

:S oh

OpenStudy (anonymous):

let me show you something maybe you'll find it a little more understanding than me

OpenStudy (wolf1728):

Are we supposed to be comparing the area of non-existant rectangles?

OpenStudy (anonymous):

i do not think so

OpenStudy (anonymous):

take a look at it please

OpenStudy (wolf1728):

Sure will

OpenStudy (anonymous):

ok thanks

OpenStudy (wolf1728):

I looked. 16 -x² was factored incorrectly. It equals (4 +x)(4 -x) They had an x-4 in their answer which is wrong.

OpenStudy (anonymous):

I'm glad you saw that because i dont know how to make sense of this

OpenStudy (wolf1728):

I was hoping they had an answer.

OpenStudy (anonymous):

me too okay so this is something someone said

OpenStudy (anonymous):

A/B (16 - x^2) / (x^2 + 2x - 24) (4 - x)(4 + x) / ( (x + 6)(x - 4) ) -(x - 4)(4 + x) / ( (x + 6)(x - 4) ) -(4 + x) / ( (x + 6) ) <===== the ratio

OpenStudy (wolf1728):

I calculated (x +6) • (x -4) / (4 +x) • (4 -x) I think they messed around with the (4-x) term. (again) How about the ratio of the areas is (x +6) • (x -4) / (4 +x) • (4 -x) ?

OpenStudy (anonymous):

?

OpenStudy (wolf1728):

or the way they posted their ratio (4 +x) • (4 -x) / (x +6) • (x -4)

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

A/B = (16 - x^2)/(x^2 + 2x -24) = ((4 - x)(4 + x))/((x + 6)(x - 4)) = (-x - 4)/(x + 6)

OpenStudy (anonymous):

what about that

OpenStudy (wolf1728):

You change (4-x) on the second line to (-x-4) on the third line and what cancels out with the x-4 in the denominator?

OpenStudy (anonymous):

I didnt do it

OpenStudy (anonymous):

:( i was just showing you something i saw

OpenStudy (wolf1728):

Sorry well somebody did it and they did it wrong. Hmmmm and no closer to the solution. Is this problem for an online test or school?

OpenStudy (anonymous):

online test flvs to be exact

OpenStudy (wolf1728):

Is there anyway you could ask a question about this problem?

OpenStudy (anonymous):

i wish i could get out the test and continue tmr but i cant it is one access only then your locked out

OpenStudy (anonymous):

My teacher is asleep i think nd she wont be back till 8 or later by that time ill be timed out the test

OpenStudy (wolf1728):

Well then I don't know what to say. Maybe the answer is (x +6) • (x -4) / (4 +x) • (4 -x)

OpenStudy (anonymous):

Okay i will take it and run with it but can you show me what to right in my answer

OpenStudy (anonymous):

then i have the last question i'll put it in a different thread afterwards

OpenStudy (wolf1728):

What I would say is the ratio of Area B to Area A is (x+6)•(x-4) / (4+x)•(4-x) This cannot be reduced any further.

OpenStudy (anonymous):

k thank you

OpenStudy (wolf1728):

hope it helps

OpenStudy (anonymous):

it did

OpenStudy (wolf1728):

Well thanks

OpenStudy (anonymous):

Yvw

OpenStudy (wolf1728):

Well maybe it's time to leave OpenStudy

OpenStudy (anonymous):

No please dont go just yet i have one more question im dying to get this test done

OpenStudy (wolf1728):

Okay - glad I just didn't leave

OpenStudy (anonymous):

Phew* new thread okay

OpenStudy (wolf1728):

Sure

OpenStudy (anonymous):

i'll tag you